# SPAR: Self-supervised Placement-Aware Representation Learning for Distributed Sensing

### v3

@@ Abstract
We present SPAR, a framework for self-supervised placement-aware representation learning in distributed sensing. Distributed sensing spans applications where multiple spatially distributed and multimodal sensors jointly observe an environment, from vehicle monitoring to human activity recognition and earthquake localization. A central challenge shared by this wide spectrum of applications, is that observed signals are inseparably shaped by sensor placements, including their spatial locations and structural roles. However, existing pretraining methods remain largely placement-agnostic. SPAR addresses this gap through a unifying principle: the duality between signals and positions. Guided by this principle, SPAR introduces spatial and structural positional embeddings together with dual reconstruction objectives, explicitly modeling how observing positions and observed signals shape each other. Placement is thus treated not as auxiliary metadata but as intrinsic to representation learning. SPAR is theoretically supported by analyses from information theory and occlusion-invariant learning. Extensive experiments on three real-world datasets show that SPAR achieves superior robustness and generalization across various modalities, placements, and downstream tasks.
@@ 1 Introduction
This paper advances the state of the art in self-supervised placement-aware representation learning, motivated by the broad class of applications we term distributed sensing. By distributed sensing, we refer to systems where multiple spatially distributed sensing points—potentially spanning diverse modalities—jointly observe an environment. This definition unifies a wide spectrum of domains, including seismic and acoustic monitoring for security (Li et al., 2025; California Institute of Technology (1926), Caltech), human activity recognition with body-worn sensors (Gu et al., 2021; Sztyler & Stuckenschmidt, 2016), vehicle monitoring in urban spaces (Bathla et al., 2022), environmental monitoring (Ullo & Sinha, 2020), and smart cities (Syed et al., 2021). These scenarios, though superficially distinct, share the common challenge of reconstructing or representing an environment from heterogeneous, distributed vantage points.
Sensor placement lies at the core of distributed sensing. A sensor’s vantage point is determined by both its spatial location (e.g., GPS coordinates of a seismic station dictating which parts of the crust it samples) and its structural role (e.g., the body location of an IMU sensor that shapes its motion patterns). Robust representation learning in this setting requires models that not only capture signal content but also interpret how those signals are mediated by spatial and structural placement.
Despite rapid progress in sensing pretraining, current approaches—whether contrastive (Ouyang et al., 2024), generative reconstruction (Kara et al., 2024b), or language-model-based (Ouyang & Srivastava, 2024)—remain largely placement-agnostic, overlooking the fact that distributed sensing signals are inseparably shaped by sensor placement. This omission limits generalization across layouts, scales, and tasks.
@@
Building on this principle, SPAR introduces three key components: (1) spatial positional embeddings encoding sensor locations, (2) structural positional embeddings capturing node-specific characteristics, and (3) dual reconstruction objectives that enforce the mutual recoverability of placements and signals with contextual awareness. Together, these elements yield a cohesive, placement-aware pretraining strategy that is broadly applicable across sensing modalities and layouts. An overview of the SPAR workflow is illustrated in Figure 1. To our knowledge, this is the first work to treat placement as a universal inductive bias for distributed sensing systems as a whole, rather than as an application-specific add-on.
We further provide theoretical analyses grounded in information theory and occlusion-invariant representation learning (Kong & Zhang, 2023) to elucidate the rationale behind our design. Experiments on three real-world datasets—covering vehicle monitoring (Li et al., 2025), human activity recognition (Sztyler & Stuckenschmidt, 2016), and earthquake localization (California Institute of Technology (1926), Caltech)—demonstrate that SPAR consistently outperforms existing approaches across diverse sensing modalities, spatial configurations, and downstream tasks.
In summary, this paper makes the following contributions: (1) We introduce SPAR, a novel, general pretraining framework for distributed sensing that explicitly models spatial layouts and node-specific characteristics, guided by the duality between positions and signals. (2) We provide theoretical analyses from information-theoretic and occlusion-invariant perspectives that explain the effectiveness of our design. (3) We validate SPAR through extensive experiments on three real-world distributed sensing datasets, demonstrating superior generalizability and robustness compared to prior methods.
@@ 2 Related Work
Pretraining and Foundation Models for Sensing. Pretraining for sensing aims to learn transferable representations from unlabeled data, enabling scalable downstream learning. Existing approaches largely fall into three paradigms: contrastive learning, generative (masked reconstruction), and LLM-based frameworks. Contrastive methods align multi-modal embeddings in a shared space. Early works such as Cosmo (Ouyang et al., 2022), Cocoa (Deldari et al., 2022), and FOCAL (Liu et al., 2023) focus on intra-sample contrast via modality-specific augmentations, while more recent models like ImageBind (Girdhar et al., 2023) and MMBind (Ouyang et al., 2024) extend to loosely paired or unpaired modalities. Generative approaches rely on masked reconstruction (Woo et al., 2024; Das et al., 2024). Ti-MAE (Li et al., 2023), MOMENT (Goswami et al., 2024), and TS-MAE (Liu et al., 2025) adapt autoencoding to time series, with TS-MAE using a continuous-time formulation. Other methods, such as FreqMAE (Kara et al., 2024b) and PhyMask (Kara et al., 2024a), introduce frequency-domain masking tailored to sensing signals. LLM-based frameworks integrate sensor data into language-centric systems (Gruver et al., 2023; Garza et al., 2023). LIMU-BERT (Xu et al., 2021) adapts masked language modeling for inertial data, while Penetrative AI (Xu et al., 2024), LLMSense (Ouyang & Srivastava, 2024), and IoT-LM (Mo et al., 2024) introduce prompting, summarization, and modality-specific adapters for cross-task transfer and zero-shot inference. While effective, these methods do not explicitly account for the spatial layout and node-specific characteristics critical to distributed sensing. In contrast, SPAR incorporates spatial and structural information directly into pretraining, enhancing contextual grounding and robustness.
@@
Pretraining with Different Notions of "Spatial" Context. Several works incorporate spatial context, though definitions of “spatial” differ. In vision, it typically refers to grids of pixels or patches, as in video (Feichtenhofer et al., 2022; Wu et al., 2023a), remote sensing (Lin et al., 2023; Reed et al., 2023; Irvin et al., 2023), and 3D medical imaging (Gu et al., 2024). Beyond vision, "spatial" often denotes discrete symbolic entities, e.g., joints in SkeletonMAE (Wu et al., 2023b), EEG channels in MV-SSTMA (Li et al., 2022a) and MMM (Yi et al., 2023), or sensor identities in Gao et al. (Gao et al., 2023) and Miao et al. (Miao et al., 2024). By contrast, our method integrates continuous node coordinates into pretraining, enabling modeling of arbitrary sensor layouts that depart from token sequences or regular grids, and generalization to unseen configurations. A related but distinct line is scene reconstruction and novel view synthesis (Mildenhall et al., 2021; Kerbl et al., 2023; Wu et al., 2024), which also exploits spatial layouts but targets synthesis quality, rather than transferable representations for sensing.
@@
Pretraining via Positional Reconstruction Objectives. A third line of work, often without using the term “spatial,” incorporates positional reconstruction objectives. In vision, Doersch et al. (Doersch et al., 2015) proposed predicting relative patch positions, extended by jigsaw (Noroozi & Favaro, 2016) and content restoration (Kim et al., 2018). DeepPermNet (Santa Cruz et al., 2017) learns permutation structures, MP3 (Zhai et al., 2022) predicts absolute patch locations, and LOCA (Caron et al., 2024) predicts relative positions of clustered patches. In NLP, StructBERT (Wang et al., 2019), ALBERT (Lan et al., 2019), and SLM (Lee et al., 2020) use sentence order prediction and sequence restoration, while Nandy et al. (Nandy et al., 2024) extend to permutation-based objectives. Beyond vision and language, GeoMAE (Tian et al., 2023) reconstructs geometric features of masked point clouds, and LEGO (Sun et al., 2024) recovers perturbed molecular geometries. While conceptually related, these methods operate on discrete, domain-specific positional targets (e.g., patch indices, sentence order). By contrast, our approach reconstructs continuous physical positions of sensor nodes, naturally aligning with distributed sensing.
@@ 3 Method
To develop a pretraining method that effectively utilizes the unique placement characteristics of sensing nodes, we propose SPAR, which explicitly leverages the duality between observer placement and signals in the distributed sensing data. Specifically, we extend the traditional MAE framework by introducing spatial positional embeddings to represent device locations, and structural positional embeddings to encapsulate effects of other placement characteristics (such as orientation). Furthermore, we propose to optimize our model with novel dual reconstruction objectives to enhance its ability to retain both signal and spatial information in its learned representations. The overall architecture of SPAR is shown in Figure 2, with each component detailed below. Importantly, SPAR is grounded in solid theoretical foundations from both information theory and the study of occlusion-invariant representations, offering deep insights into our design.
@@ 3.1 Embedding for Signals, Spatial Positions, and Structural Positions
We consider a multi-modality distributed sensing system with modalities, where the -th modality () consists of sensing nodes. The signals collected from these nodes are first tokenized to be compatible with transformer encoders. The tokenization strategy is task-specific and flexible—for example, IMU time-series data can be divided into temporal segments, while acoustic spectrograms can be split into patches. We denote the tokenized signals as , where is the number of tokens and is the token dimension. We then project these tokens into the transformer embedding space using a learnable linear layer, as , yielding signal embeddings , where is the transformer model dimension.
A distinguishing aspect of distributed sensing data is the availability of spatial positions of the nodes, reflecting their physical layout in the field, which can be denoted as . For instance, in Figure 2, the spatial positions are two-dimensional, indicating longitudinal and lateral node locations. Unlike the discrete ordinal indices typically used in NLP (Vaswani et al., 2017) or CV (Dosovitskiy et al., 2020), spatial positions in distributed sensing data are continuous vectors, making classical positional embedding strategies unsuitable (Vaswani et al., 2017; Dosovitskiy et al., 2020; Su et al., 2024; Press et al., 2021). To address this, we propose to continuously project the spatial positions into the embedding space as , where is the spatial positions broadcasted to match the dimension of the tokenized signals. The spatial positional embeddings are then added to the signal embeddings to incorporate spatial context into the model.
@@
Beyond this learnable formulation, we further explore leveraging Large Language Models (LLMs) to derive structural positional embeddings from free-form textual metadata describing each sensor’s placement, modality, and signal characteristics. These LLM-derived embeddings, obtained by encoding textual descriptions, are projected and fused with the other embeddings but kept frozen during pretraining. This metadata-driven variant, which we denote as SPAR+LLM, encourages generalization to previously unseen sensors and placements.
Structural positions bear an interesting mathematical interpretation: if we assume the influence of a node’s unique characteristics on its signal embedding can be summarized as an additive vector, which lies within a specific subspace of the embedding space, then the weight matrix of can be understood as a learned set of basis vectors spanning this subspace. The structural position of each node can then be viewed as the coordinate of the corresponding additive vector in this subspace, thereby substantiating its meaning as an abstract “position”.
@@ 3.3 Theoretical Analyses
As shown by Kong et al. (Kong & Zhang, 2023), classical MAE can be viewed as a form of contrastive learning, where the positive pair consists of two complementary masked views of the signals:
@@ 4 Evaluation
In this section, we present our experimental evaluation of SPAR on three multiple multi-modal, multi-node distributed sensing datasets spanning diverse sensing modalities and spatial scales. To ensure a fair comparison, all baseline methods and our model use the same ViT backbone architecture (Dosovitskiy et al., 2020) and identical task-specific prediction heads. Pretraining and fine-tuning are conducted for the same number of epochs across all methods. All reported results are aggregated over three random seeds. The prediction heads are designed to be lightweight and straightforward, tailored to the needs of each downstream task. Detailed descriptions of each task setup can be found in Appendix E.
Datasets. We conducted experiments on three real-world distributed sensing datasets: (1) the M3N-VC dataset(Li et al., 2025), which includes acoustic and seismic signals from moving vehicles, collected across six distinct outdoor scenes; (2) the Ridgecrest Seismicity Dataset(California Institute of Technology (1926), Caltech), containing multi-modal seismic waveform recordings of earthquake events in the Ridgecrest region of California; and (3) the RealWorld-HAR dataset(Sztyler & Stuckenschmidt, 2016), comprising accelerometer, gyroscope, and magnetometer readings for human activity recognition. Further dataset details are available in Appendix E.
Baselines. We compare SPAR against eight state-of-the-art baseline methods: CMC (Tian et al., 2020), Cosmo (Ouyang et al., 2022), SimCLR (Chen et al., 2020), AudioMAE (Huang et al., 2022), CAV-MAE (Gong et al., 2022), FOCAL (Liu et al., 2023), FreqMAE (Kara et al., 2024b), and PhyMask (Kara et al., 2024a). Among these, CMC, Cosmo, SimCLR, and FOCAL are contrastive learning-based methods, while AudioMAE, CAV-MAE, and FreqMAE follow the masked autoencoding (MAE) paradigm. Please see Appendix E.1 for a detailed description of each baseline.
@@ 4.1 Evaluation on M3N-VC Dataset
Method M3N-VC Single-vehicle Localization Label Ratio 1.0 Label Ratio 0.5 Label Ratio 0.2 MSE () () Dist. Err. () () MSE () () Dist. Err. () () MSE () () Dist. Err. () () CMC 51.11 14.67 6.76 0.75 71.81 15.32 7.99 0.64 111.37 8.02 11.05 0.57 Cosmo 38.40 4.14 6.03 0.21 53.12 9.75 7.19 0.40 97.08 9.49 10.95 0.57 SimCLR 34.40 4.47 5.64 0.25 45.14 7.34 6.57 0.08 74.53 3.13 9.48 0.17 AudioMAE 22.36 0.49 5.40 0.11 30.12 2.97 6.33 0.28 41.75 3.30 7.47 0.28 CAV-MAE 18.85 0.41 5.06 0.04 22.90 0.82 5.58 0.12 24.84 0.33 5.78 0.10 FOCAL 32.43 4.68 5.37 0.22 40.84 2.82 6.20 0.19 69.62 5.62 8.50 0.35 FreqMAE 29.61 2.85 5.36 0.16 42.06 14.44 6.25 0.70 91.40 35.32 9.15 1.27 PhyMask 28.02 5.91 5.29 0.33 33.74 2.18 5.85 0.12 64.36 4.70 8.44 0.36 SPAR 12.98 0.11 4.20 0.07 15.07 1.03 4.51 0.09 21.36 0.62 5.40 0.04
Method M3N-VC Multi-vehicle Joint Classification and Localization mAP@4m (%) () mAP@6m (%) () mAP@8m (%) () mAP@10m (%) () CMC 0.06 0.05 0.48 0.36 1.61 1.10 3.62 2.19 Cosmo 0.16 0.05 1.66 0.23 4.77 0.72 9.52 1.20 SimCLR 0.31 0.14 2.22 0.58 6.53 1.24 13.07 2.08 AudioMAE 1.39 0.48 6.96 1.42 17.11 3.24 28.98 4.01 CAV-MAE 22.12 2.94 52.08 4.16 73.41 3.24 85.36 1.78 FOCAL 0.08 0.05 0.82 0.40 2.94 1.04 6.82 1.99 FreqMAE 0.24 0.01 1.67 0.32 5.34 0.99 11.31 1.49 PhyMask 0.08 0.03 0.88 0.24 3.04 0.74 6.64 1.46 SPAR 41.57 2.69 71.82 3.69 86.28 1.77 92.99 0.79
Method M3N-VC Classification Ridgecrest Earthquake Localization RealWorld-HAR Recognition Accuracy (%) () F1 (%) () MSE () () Dist. Err. () () Accuracy (%) () F1 (%) () CMC 89.53 7.62 89.33 7.78 94.25 6.67 10.38 0.63 74.97 1.23 74.82 2.18 Cosmo 94.21 0.50 94.04 0.54 98.24 13.77 10.44 0.83 84.37 0.33 85.30 0.43 SimCLR 95.53 0.73 95.41 0.74 99.87 11.31 10.29 0.52 84.36 0.47 85.49 0.36 AudioMAE 99.06 0.23 99.03 0.24 33.65 3.51 5.65 0.29 89.18 0.32 90.11 0.53 CAV-MAE 98.97 0.04 98.94 0.04 31.58 3.57 5.48 0.37 88.12 0.24 89.05 0.35 FOCAL 93.62 0.75 93.46 0.76 131.50 1.48 12.53 0.09 84.98 0.73 86.24 0.77 FreqMAE 92.72 0.75 92.55 0.79 54.08 5.44 7.14 0.25 83.43 0.56 84.07 0.50 PhyMask 83.38 2.33 82.68 2.27 56.39 3.27 7.67 0.39 84.79 3.23 82.15 9.13 SPAR 99.27 0.07 99.26 0.07 23.46 2.77 5.37 0.24 89.63 0.57 90.45 0.63
@@
We begin with the M3N-VC dataset, focusing on the task of single-vehicle localization, where the goal is to predict the position of a vehicle within the monitored area. We pretrain on the full dataset and finetune only the prediction head (a single transformer layer) on scene "H24," which contains a single moving vehicle. To test robustness under limited supervision, we vary the ratio of labeled data from 100% to 20%. As shown in Table 1, SPAR consistently achieves the lowest MSE and Distance Error across all label ratios, demonstrating resilience to scarce supervision. Example localization visualizations are provided in Figure 3(a) (Appendix B).
@@
The second task is single-vehicle classification, where the model distinguishes among four vehicle types and a background class. The setup mirrors that of localization. As reported in Table 3, SPAR attains the highest accuracy and F1 score among all methods. The confusion matrix (Figure 5) and T-SNE plot (Figure 4) confirm that the learned embeddings cleanly separate all five classes, underscoring their discriminative power.
@@
We next consider the more challenging task of multi-vehicle joint classification and localization, where multiple vehicles move simultaneously, generating overlapping signals. We pretrain on the full dataset and finetune on scene "I22," which includes multiple vehicles. A 2-layer transformer head with a DETR-style loss (Carion et al., 2020) is used. To evaluate performance, we adopt mAP@r from object detection (Lin et al., 2014), where predictions are correct only if both class and location are accurate within radius r. As shown in Table 2, SPAR significantly outperforms all baselines across thresholds, including strict ones (e.g., 4m), despite noisy 1Hz smartphone GPS labels. This highlights its strong spatial reasoning under complex conditions. Additional visualizations of predictions are included in Figure 6 (Appendix B).
Finally, we conduct three complementary evaluations (details in Appendix A): (1) Lossy Communication: SPAR remains robust under random node-level data dropout, outperforming baselines (Table 6). (2) Unseen Sensor Placements: The pretrained, frozen encoders of SPAR generalize well to unseen placement, confirming placement-aware robustness (Table 7). (3) Ablations: The ablation studies indicate that each design component in SPAR contributes meaningfully and synergically, and SPAR maintains robust to varing hyperparameters (Table 8).

### v4

@@ Abstract
We present SPAR, a framework for self-supervised placement-aware representation learning in distributed sensing. Distributed sensing spans applications where multiple spatially distributed and multimodal sensors jointly observe an environment, from vehicle monitoring to human activity recognition and earthquake localization. A central challenge shared by this wide spectrum of applications is that observed signals are inseparably shaped by sensor placements, including their spatial locations and structural characteristics. However, existing pretraining methods remain largely placement-agnostic. SPAR addresses this gap through a unifying principle: the duality between signals and positions. Guided by this principle, SPAR introduces spatial and structural positional embeddings together with dual reconstruction objectives, explicitly modeling how observing positions and observed signals shape each other. Placement is thus treated not as auxiliary metadata but as intrinsic to representation learning. SPAR is theoretically supported by analyses from information theory and occlusion-invariant learning. Extensive experiments on three real-world datasets show that SPAR achieves superior robustness and generalization across various modalities, placements, and downstream tasks.
@@ 1 Introduction
This paper advances self-supervised placement-aware representation learning, motivated by the broad class of applications we term distributed sensing. By distributed sensing, we refer to systems where multiple spatially distributed sensing points, potentially spanning diverse modalities, jointly observe an environment. This definition unifies a wide spectrum of domains, including seismic and acoustic monitoring for security (li2025restoreml; caltech1926socal), human activity recognition with body-worn sensors (gu2021survey; sztyler2016body), vehicle monitoring in urban spaces (bathla2022autonomous), environmental monitoring (ullo2020advances), and smart cities (syed2021iot). These scenarios, though superficially distinct, share the common challenge of reconstructing or representing an environment from heterogeneous, distributed vantage points.
Sensor placement lies at the core of distributed sensing. A sensor’s vantage point is determined by both its spatial location (e.g., GPS coordinates of roadside microphones monitoring passing vehicles, governing the propagation paths of acoustic waves) and its structural characteristic (e.g., differences in sensor mounting, orientation, or coupling to the ground, which shape frequency response and noise patterns). Robust representation learning in this setting requires models that not only capture signal content but also interpret how those signals are mediated by sensor placement.
Despite rapid progress in pretraining, current approaches, whether contrastive (ouyang2024mmbind), generative reconstruction (kara2024freqmae), or language-model-based (ouyang2024llmsense), remain largely placement-agnostic, overlooking the fact that distributed sensing signals are inseparably shaped by sensor placement. This omission limits generalization across layouts and tasks.
@@
Building on this principle, SPAR introduces explicit spatial and structural positional embeddings, and introduces dual reconstruction objectives that enforce the mutual recoverability of placements and signals with contextual awareness. Together, these elements yield a cohesive, placement-aware pretraining strategy that is broadly applicable across sensing modalities and layouts. An overview of the SPAR workflow is illustrated in Figure 1. To our knowledge, this is the first work to treat placement as a universal inductive bias for distributed sensing systems as a whole, rather than as an application-specific add-on.
We further provide theoretical analyses grounded in information theory and occlusion-invariant representation learning (kong2023understanding) to elucidate the rationale behind our design. Experiments on three real-world datasets, covering vehicle monitoring (li2025restoreml), human activity recognition (sztyler2016body), and earthquake localization (caltech1926socal), demonstrate that SPAR consistently outperforms existing approaches across diverse sensing modalities, spatial configurations, and downstream tasks.
In summary, this paper makes the following contributions: (1) We introduce SPAR, a novel, general pretraining framework for distributed sensing that explicitly models spatial layouts and node-specific characteristics, guided by the duality between positions and signals. (2) We provide theoretical analyses from information-theoretic and occlusion-invariant perspectives that explain the effectiveness of our design. (3) We validate SPAR through extensive experiments on three real-world distributed sensing datasets, demonstrating superior generalizability and robustness compared to prior methods. We release our implementation and scripts at https://anonymous.4open.science/r/SPAR-F6F3/.
@@ 2 Related Work
Pretraining and Foundation Models for Sensing. Pretraining for sensing seeks transferable representations from unlabeled data and broadly follows three paradigms: contrastive learning, masked reconstruction, and LLM-based frameworks. Contrastive methods align multi-modal embeddings, ranging from intra-sample approaches such as Cosmo (ouyang2022cosmo), Cocoa (deldari2022cocoa), and FOCAL (liu2023focal), to large-scale or loosely paired models like ImageBind (girdhar2023imagebind) and MMBind (ouyang2024mmbind). Generative approaches are dominated by masked autoencoding (woo2024unified; das2024decoder), including time-series adaptations such as Ti-MAE (li2023ti), MOMENT (goswami2024moment), TS-MAE (liu2025ts), and frequency-aware variants like FreqMAE (kara2024freqmae) and PhyMask (kara2024phymask). LLM-based frameworks integrate sensing into language-centric models (gruver2023large; garza2023timegpt), including LIMU-BERT (xu2021limu), Penetrative AI (xu2024penetrative), LLMSense (ouyang2024llmsense), and IoT-LM (mo2024iot). Despite their success, these methods remain largely placement-agnostic, overlooking the spatial layout and node-specific characteristics intrinsic to distributed sensing; in contrast, SPAR explicitly incorporates spatial and structural information during pretraining to improve contextual grounding and robustness.
@@
Pretraining with Different Notions of "Spatial" Context. Several works incorporate spatial context, though definitions of “spatial” differ. In vision, it typically refers to grids of pixels or patches, as in video (feichtenhofer2022masked; wu2023dropmae), remote sensing (lin2023ss; reed2023scale; irvin2023usat), and 3D medical imaging (gu2024self). Beyond vision, "spatial" often denotes discrete symbolic entities, e.g., joints in SkeletonMAE (wu2023skeletonmae), EEG channels in MV-SSTMA (li2022multi) and MMM (yi2023learning), or sensor identities in Gao et al. (gao2023spatial) and Miao et al. (miao2024spatial). By contrast, our method integrates continuous node coordinates into pretraining, enabling modeling of arbitrary sensor layouts. Another related line is scene reconstruction and novel view synthesis (mildenhall2021nerf; kerbl20233d; wu20244d), which also exploits spatial layouts but targets synthesis quality, rather than transferable representations for sensing.
@@
Pretraining via Positional Reconstruction Objectives. A third line of work, often without using the term “spatial,” incorporates positional reconstruction objectives. In vision, Doersch et al. (doersch2015unsupervised) proposed predicting relative patch positions, extended by jigsaw (noroozi2016unsupervised) and content restoration (kim2018learning). DeepPermNet (santa2017deeppermnet) learns permutation structures, MP3 (zhai2022position) predicts absolute patch locations, and LOCA (caron2024location) predicts relative positions of clustered patches. In NLP, StructBERT (wang2019structbert), ALBERT (lan2019albert), and SLM (lee2020slm) use sentence order prediction and sequence restoration, while Nandy et al. (nandy2024order) extend to permutation-based objectives. Beyond vision and language, GeoMAE (tian2023geomae) reconstructs geometric features of masked point clouds, and LEGO (sun20243d) recovers perturbed molecular geometries. While conceptually related, these methods operate on discrete, domain-specific positional targets (e.g., patch indices, sentence order). By contrast, our approach reconstructs continuous physical positions of nodes in distributed sensing applications.
@@ 3 Method
To develop a pretraining method that effectively utilizes the unique placement characteristics of sensing nodes, we propose SPAR, which explicitly leverages the duality between observer placement and signals in the distributed sensing data. Specifically, we extend the traditional MAE framework by introducing explicit spatial and structural positional embeddings, to represent device locations and to encapsulate effects of other placement characteristics (such as orientation). Furthermore, we propose to optimize our model with dual reconstruction objectives to enhance its ability to retain both signal and spatial information in its learned representations. The overall architecture of SPAR is shown in Figure 2, with each component detailed below. Importantly, SPAR is grounded in solid theoretical foundations from both information theory and the study of occlusion-invariant representations, offering deep insights into our design.
@@ 3.1 Embedding for Signals, Spatial Positions, and Structural Positions
We consider a multi-modality distributed sensing system with modalities, where the -th modality () consists of sensing nodes. The signals collected from these nodes are first tokenized to be compatible with transformer encoders. The tokenization strategy is task-specific and flexible. For example, IMU time-series data can be divided into temporal segments, while acoustic spectrograms can be split into patches. We denote the tokenized signals as , where is the number of tokens and is the token dimension. We then project these tokens into the transformer embedding space using a learnable linear layer, as , yielding signal embeddings , where is the transformer model dimension.
A distinguishing aspect of distributed sensing data is the availability of spatial positions of the nodes, reflecting their physical layout in the field, which can be denoted as . For instance, in Figure 2, the spatial positions are two-dimensional, indicating longitudinal and lateral node locations. Unlike the discrete ordinal indices typically used in NLP (vaswani2017attention) or CV (dosovitskiy2020image), spatial positions in distributed sensing data are continuous vectors, making classical positional embedding strategies unsuitable (vaswani2017attention; dosovitskiy2020image; su2024roformer; press2021train). To address this, we propose to continuously project the spatial positions into the embedding space as , where is the spatial positions broadcasted to match the dimension of the tokenized signals. The spatial positional embeddings are then added to the signal embeddings to incorporate spatial context into the model.
@@ 3.3 Theoretical Analyses
As shown by Kong et al. (kong2023understanding), classical MAE can be viewed as a form of contrastive learning, where the positive pair consists of two complementary masked views of the signals:
@@ 4 Evaluation
In this section, we present our experimental evaluation of SPAR on three multiple multi-modal, multi-node distributed sensing datasets spanning diverse sensing modalities and spatial scales. To ensure a fair comparison, all baseline methods and our model use the same ViT backbone architecture (dosovitskiy2020image) and identical task-specific prediction heads. Pretraining and fine-tuning are conducted for the same number of epochs across all methods. All reported results are aggregated over three random seeds. The prediction heads are designed to be lightweight and straightforward, tailored to the needs of each downstream task. Detailed descriptions of each task setup can be found in Appendix E.
Datasets. We conducted experiments on three real-world distributed sensing datasets: (1) the M3N-VC dataset(li2025restoreml), which includes acoustic and seismic signals from moving vehicles, collected across six distinct outdoor scenes; (2) the Ridgecrest Seismicity Dataset(caltech1926socal), containing multi-modal seismic waveform recordings of earthquake events in the Ridgecrest region of California; and (3) the RealWorld-HAR dataset(sztyler2016body), comprising accelerometer, gyroscope, and magnetometer readings for human activity recognition. Further dataset details are available in Appendix E.
Baselines. We compare SPAR against eight baseline methods: CMC (tian2020contrastive), Cosmo (ouyang2022cosmo), SimCLR (chen2020simple), AudioMAE (huang2022masked), CAV-MAE (gong2022contrastive), FOCAL (liu2023focal), FreqMAE (kara2024freqmae), and PhyMask (kara2024phymask). These baselines are chosen to provide (i) paradigm coverage across contrastive learning (CMC, Cosmo, SimCLR, FOCAL) and masked autoencoding (AudioMAE, CAV-MAE, FreqMAE), and (ii) domain coverage spanning widely used, strong general-purpose pretraining methods for time-series/multimodal representation learning (e.g., AudioMAE, CAV-MAE, CMC, SimCLR) as well as recent sensing-tailored approaches (e.g., Cosmo, FOCAL, FreqMAE, PhyMask). Please see Appendix E.1 for detailed descriptions.
@@ 4.1 Evaluation on M3N-VC Dataset
[…]
Method Single-vehicle Localization MSE () Dist. Err. () SPAR 12.98 0.11 4.20 0.07 w/o Geometric Augmentation in Pretrain 15.59 0.56 4.67 0.04 w/o Reconstructing Spatial Positions 14.73 0.35 4.62 0.03 w/o Spatial Positional Embedding 15.12 0.58 4.67 0.07 w/o Structural Positional Embedding 22.55 2.98 5.08 0.13 + Node-Drop Masking 17.71 3.17 4.82 0.33 + Node-Balanced Masking 12.54 1.63 4.14 0.21 + Mask Ratio 0.85 13.89 2.17 4.35 0.24 + Mask Ratio 0.5 14.81 2.57 4.45 0.31
[…]
@@
Single-vehicle Localization. We begin with the M3N-VC dataset, focusing on the task of single-vehicle localization, where the goal is to predict the position of a vehicle within the monitored area. We pretrain on the full dataset and finetune only the prediction head (a single transformer layer) on scene "H24," which contains a single moving vehicle. To test robustness under limited supervision, we vary the ratio of labeled data from 100% to 20%. As shown in Table 1, SPAR consistently achieves the lowest MSE and Distance Error across all label ratios, demonstrating resilience to scarce supervision. Example localization visualizations are provided in Figure 4(a) (Appendix B).
@@
Single-vehicle Classification. In this task, we aim to distinguish among four vehicle types and a background class. The setup mirrors that of localization. As reported in Table 5, SPAR attains the highest accuracy and F1 score among all methods. The confusion matrix (Figure 5) and T-SNE plot (Figure 6) confirm that the learned latent embeddings cleanly separate all five classes, underscoring their discriminative power.
@@
Multi-vehicle Detection. We next consider a more challenging task, where we jointly localize and classify multiple vehicles moving simultaneously, which is similar to the object detection task in CV. We pretrain on the full dataset and finetune on the scene "I22," which includes multiple vehicles. A 2-layer transformer head with a DETR-style loss (carion2020end) is used as the detection head. To evaluate performance, we adopt mAP@r from object detection (lin2014microsoft), where predictions are correct only if both class and location are accurate within radius r. As shown in Table 2, SPAR significantly outperforms all baselines across thresholds, including strict ones (e.g., 4m), despite noisy 1Hz smartphone GPS labels. This highlights its strong spatial reasoning under complex conditions. Additional visualizations of predictions are included in Figure 7 (Appendix B).
Unseen Sensor Placements. Next, we assess generalization to unseen sensor placements. All models are pretrained on the full dataset, excluding scenes H08 and H24 (which share similar configurations), then finetuned and evaluated on H24. To simulate transfer, nodes in H24 are assigned structural position vectors randomly drawn from those learned during pretraining. As reported in Table 3, SPAR continues to outperform baselines, underscoring its placement-aware generalization ability.
Ablation Study. Finally, we conduct ablations to quantify the contribution of each design choice (Table 4). Removing geometric augmentation, the spatial reconstruction objective, or spatial embeddings all lead to comparable performance drops, highlighting their complementary roles. Excluding structural embeddings results in the most severe degradation, underscoring their critical role in modeling node-specific characteristics. We also test alternative masking strategies: Node-Drop Masking (masking entire nodes) reduces performance, while Node-Balanced Masking (ensuring a minimum number of unmasked tokens per node) offers slight gains over random masking. We further vary the mask ratio (0.85 and 0.5) and observe only minor performance changes relative to the default 0.75, indicating the robustness of the framework.
Finally, we conduct another three complementary evaluations that are detailed in Appendix A: (1) Robustness to Lossy Communication: SPAR remains robust under random node-level data dropout, outperforming baselines (Table 7). (2) Preservation of Spatial Information in Learned Representations: We provide evidence that SPAR preserves spatial information in its representations via a spatial reconstruction ablation study (Figure 3(a)). (3) Robustness to Noise in Spatial Positions: SPAR’s learned representation is robust to noise in nodes’ spatial positions.